For the following problem, I know A = l w, is it l=w+2 and A=60? 60=w+2(w) being the final?
The length of a rectangle is 2 cm greater than twice its width. If the area of the rectangle is 60 cm , find the length of the rectangle.

Answer by stanbon(48558)

The length of a rectangle is 2 cm greater than twice its width. If the area of the rectangle is 60 cm , find the length of the rectangle.
-------------
Length = 2width + 2 or 2w+2: correct
Then Area = w(x+2): you need the parenthesis.
w(2w+2)=60
2w^2+2w-60=0
w=[-2+-sqrt(4-4(2)(-60))]/4
w=[-2+sqrt484]/4
w=[(-2+22]/4
w=5
length = 2w+2=12
Cheers,
Stan H.